OpenStudy (anonymous):

how do u solve by factoring, completing the sqaure or using the quadratic formula on this problem x^2-8x=5

OpenStudy (anonymous):

x^(2-8x) ?

OpenStudy (dumbcow):

For factoring set it equal to zero ->x^2 -8x -5 = 0 Then see if any factors of -5 add up to -8 if there are none it is prime For completing the square, take half of -8 which is -4 and square it subtract that from -5 _> x^2-8x-5 = (x-4)^2-5-(-4)^2 = (x-4)^2 -21 = 0 For quadratic formula which really is the same as completing the square use the formula x = -b +-sqrt(b^2-4ac)/2a

OpenStudy (anonymous):

k thank you soooo much i used the completeting the sqaure. so thats the answer

OpenStudy (dumbcow):

once you solve for x yes (x-4)^2 = 21 x-4 = sqrt(21) x= 4 +- sqrt(21)